OpenStudy (anonymous):

hey @midhun.madhu1987 help me give me some important identites that uwd help me prove l.hs is eqaul to rhs in trignometry

OpenStudy (midhun.madhu1987):

Question please...

OpenStudy (midhun.madhu1987):

:0

OpenStudy (anonymous):

wuttt

OpenStudy (midhun.madhu1987):

ok

OpenStudy (anonymous):

thnxxx

OpenStudy (anonymous):

u deree

OpenStudy (midhun.madhu1987):

\[\sin ^{2}\theta+\cos ^{2}\theta =1\] \[\sec ^{2}\theta=1+tan ^{2}\theta\] \[cosec ^{2}\theta=1+cot ^{2}\theta\]

OpenStudy (anonymous):

aything ode rthan theeeess

OpenStudy (midhun.madhu1987):